sin theta + sin phi = a cos theta + cos phi = b theta- phi = alpha eliminate theta and phi

sin theta  + sin phi = a
cos theta + cos phi = b
theta- phi = alpha
eliminate theta and phi

Grade:11

1 Answers

mohan
6 years ago
sin(theta) +sin(phi)=a implies 2sin((theta +phi)/2)cos((theta-phi)/2)=a “equation1” similarly 2cos((theta+phi)/2)cos((theta-phi)/2)=b “equation 2” on dividing both equations we have tan((theta+phi)/2)=a/b implies sin((theta+phi)/2)=a/(a2+b2)1/2 substituting in equation  one we have cos(alpha/2)=(a2+b2)1/2/2